OpenStudy (anonymous):

The gravitational attraction A between two masses varies inversely as the square of the distance between them. The force is 2.25 lb when the masses are 4 ft apart, what is the attraction when the masses are 6 ft apart?

OpenStudy (anonymous):

2.25/4=x/6 4x=13.5 x=13.5/4lb

OpenStudy (anonymous):

hmm i got A=k/(d^2) 2.25=k/16 36=k A=36/(d^2) A=36/(6^2) A=36/36=1